OpenStudy (anonymous):

Find the vertex of this parabola: y=x^2-2x-5

OpenStudy (anonymous):

vertex = -b/2a = 2/2 = 1 when x=1 y = 1-2-5 = -6 vertex ia at (1,-6)

OpenStudy (anonymous):

\[y=(x-1)^2-5-(1*1) \[y=(x-1)^2-6\]x=1 y=-6